OpenStudy (anonymous):

When finding the sum of the arithmetic sequence 4 + 1 + –2 + –5 + –8 + –11, you will have a step of work that contains the number –3. Use complete sentences to describe how the –3 was calculated and what it means.

OpenStudy (anonymous):

-3 is the constant difference abbreviated as d as is defined as the change between the preceding term and the previous term i.e (\[s _{n+1} - s _{n}\] in your case consider 1st, 2nd and 3rd term: 3rd term - 2nd term = 2nd term - 1st term = -2-1 = 1 - 4 = -3